Find the steady state temperature of the black body disc having radius 3cm and mass 5g at 10 0C surrounding temperature,when kept initially at -10 0C.
1
Expert's answer
2020-12-15T11:33:58-0500

Temperature is given by

"T=\frac{1}{\sqrt[3]{\frac{3\pi r^2 \sigma t}{mc}}+\frac{1}{(T_s-T_{sur}) ^3}}+T_{sur}"

When t goes to infinite then

"T=T_{sur}=273+10=283K"